# Characterizing the minimax rate of nonparametric regression under bounded convex constraints

### v3

@@ 2.1 Lower Bound
We use (wainwright2019high, (15.52)), the decoupling property of KL-divergence, and the decomposition of the joint distribution of the $(X_{i},Y_{i})$. Let $\mathbb{P}_{Y_{i}|X_{i}}^{f_{j}}$ be the distribution of $(Y_{i}|X_{i},J=j)$, which is a normal distribution with mean $f_{j}(X_{i})$ and variance $\sigma^{2}$. Denote by $\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}$ the distribution of $(Y_{i},X_{i}|J=j)$ and $\mathbb{P}_{X_{i}}$ that of $X_{i}$, so that $\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}=\mathbb{P}_{X_{i}}\times\mathbb{P}_{Y_{i}|X_{i}}^{f_{j}}$. Note that the $X_{i}$ does not depend on $f_{j}$. Similarly, for any function $h\in\mathcal{F}$, if we set $Y_{i}=h(X_{i})+\xi_{i}$ instead, denote the analogous distributions by $\mathbb{P}_{(X_{i},Y_{i})}^{h}=\mathbb{P}_{X_{i}}\times\mathbb{P}_{Y_{i}|X_{i}}^{h}$. Then we obtain
$I(\vec{X},\vec{Y};J)$ $\leq\frac{1}{m}\sum_{j}D_{KL}\left(\prod_{i}\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}\parallel\prod_{i}\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)$
$=\frac{1}{m}\sum_{j}\sum_{i}D_{KL}\left(\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}\parallel\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)$
@@
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{X_{i}}\left[D_{KL}\left(N(f_{j}(X_{i}),\sigma^{2})\parallel N(h(X_{i}),\sigma^{2})\right)\right]$
@@
Lemma 2.3.
We have
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq\frac{\varepsilon^{2}}{8c^{2}},$
for any $\varepsilon$ satisfying $\log M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c)>4\left(\frac{n\varepsilon^{2}}{2\sigma^{2}}\vee\log 2\right)$, where $c$ is the constant from the Definition 1.1 which is fixed to some large enough value.
{toappendix}
@@
Pick a $f\in\mathcal{F}$ that achieves $M:=M_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c):=\sup_{f\in\mathcal{F}}M(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F})$ (otherwise, repeat the following argument by picking a limiting sequence of functions). Let $f_{1},\dots,f_{M}$ be a maximal $\varepsilon/c$-packing set for $B(f,\varepsilon)\cap\mathcal{F}$, so that $\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>\varepsilon/c$ for all $i\neq j$. Because $f_{1},\dots,f_{M}\in B(f,\varepsilon)\cap\mathcal{F}$, it follows for each $j\in[M]$ that $\|f_{j}-f\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon$. Letting $J$ be uniformly distributed on $[M]$, we have from Lemma 2.2 that
@@
$\inf_{\widehat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{X,Y}\|\widehat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $\geq\frac{\varepsilon^{2}}{4c^{2}}\bigg{(}1-\frac{\frac{n\varepsilon^{2}}{2\sigma^{2}}+\log 2}{\log M}\bigg{)}.$
@@
2.2 Upper Bound
We now propose Algorithm 1, which produces an $f^{\ast}$ achieving the minimax rate up to constants, using a procedure similar to that in neykov2022minimax. We begin by providing an informal description. We will take $c=2(C+1)$ where $C>3$ is some universal constant. In the following section, all distances are with respect to the $L_{2}(\mathbb{P}_{X})$-norm.
Pick an arbitrary $f^{\ast}\in\mathcal{F}$. Recalling that $d$ is the diameter of $\mathcal{F}$, form a maximal $\frac{d}{c}$-packing set $M_{1}$ of the set $B(f^{\ast},d)\cap\mathcal{F}$. We then update $f^{\ast}$ by reassigning it to be the least squares estimator over the $|M_{1}|$ functions in $M_{1}$:
$f^{\ast}=\mathop{\mathrm{argmin}}_{f\in M_{1}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}.$
The process now repeats but with a smaller ball and packing set (divide $d/c$ or $d$, respectively, by the next power of 2). That is, we now form a maximal $\frac{d}{2c}$-packing set $M_{2}$ of the set $B(f^{\ast},d/2)\cap\mathcal{F}$ and reassign
$f^{\ast}=\mathop{\mathrm{argmin}}_{f\in M_{2}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}.$
Then iterate the process for as many steps as desired, e.g., until some stopping condition is attained. Note that by construction of the packing sets, the updates form a Cauchy-sequence and the outputs will converge. To see this, denote by $\Upsilon_{j}$ the $j$th update of $f^{\ast}$, where $\Upsilon_{1}$ is our starting choice. Then for $j<k$, we have
$\|\Upsilon_{j}-\Upsilon_{k}\|_{L_{2}(\mathbb{P}_{X})}$ $\leq\sum_{i=j}^{k-1}\|\Upsilon_{i+1}-\Upsilon_{i}\|_{L_{2}(\mathbb{P}_{X})}\leq\sum_{i=j}^{k-1}\frac{d}{2^{i-1}}\leq\frac{d}{2^{j-2}}.$ (2.2)
Thus for any $\varepsilon>0$, for sufficiently large $N$, if $j,k\geq N$, we have $\|\Upsilon_{j}-\Upsilon_{k}\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon$. Recall moreover that $\mathcal{F}$ is complete. The packing sets may become infinite if $\mathcal{F}$ is not totally bounded. But the algorithm’s output remains a Cauchy sequence.
The construction of the packing sets is done in a data-independent manner by forming an infinite tree of possible updates of $f^{\ast}$. After forming the first packing set, the sample space $\mathcal{X}^{n}\times\mathcal{Y}^{n}$ is partitioned into $|M_{1}|$ regions corresponding to which of the $|M_{1}|$ functions is the least squares estimator. For each possible choice of updated $f^{\ast}$, a new packing set is formed, and again the sample space is partitioned. Performing the algorithm on a particular data realization corresponds to traversing a single (infinite) path down the tree.
This procedure may encounter ties in the minimization step, so we must modify our algorithm to handle such cases and in doing so, preserve measurability of the estimator it outputs. Our goal is to induce a lexicographic ordering on our functions and then choose the smallest function in the ordering to break ties. The details are given in the appendix (Remark 2.4).
{toappendix}
Remark 2.4 (Breaking Ties in Algorithm).
@@ 2.2 Upper Bound
By Theorem A.1 in gyorfi_2022, the set of continuous functions is dense in $\mathcal{F}$. Hence we can construct all of our packing sets using continuous functions only. That is, if $\mathcal{F}_{c}$ denote the set of all continuous functions in $\mathcal{F}$, then we construct $M_{k}$ as the maximal packing of $B(f^{*},d/2^{k-1})\cap\mathcal{F}_{c}$. This $M_{k}$ clearly has cardinality that is smaller than or equal to the $M_{k}$ in our previous notation since $\mathcal{F}_{c}\subseteq\mathcal{F}$. In addition since the set of all continuous functions is dense in $\mathcal{F}$, we have that for any $g\in\mathcal{F}$ there exists an $h\in\mathcal{F}_{c}$ such that $\|g-h\|\leq\varepsilon$ for any $\varepsilon>0$. Pick the point from $f\in M_{k}$ which is $d/2^{k-1}$ close to $h$ and conclude that $\|f-g\|\leq\varepsilon+d/2^{k-1}$. Let $\varepsilon$ go to zero, and we conclude that any point in $\mathcal{F}$ can be approximated by a point in $\mathcal{F}_{c}$ within distance $d/2^{k-1}$. Now suppose the domain of the functions is $\mathbb{R}^{p}$. Take the set $\mathbb{Q}^{p}$ which is dense in $\mathbb{R}^{p}$. This is a well ordered set since there is a bijection between $\mathbb{Q}^{p}$ and $\mathbb{N}$. Hence we can look at the functions from $M_{k}$ and their restrictions to $\mathbb{Q}^{p}$ and order them lexicographically. There will not be any ties, since the functions are separated and continuous (hence the functions cannot be the same on $\mathbb{Q}^{p}$ otherwise they are the same on $\mathbb{R}^{p}$ which is a contradiction) so we can take the smallest function in the lexicographic ordering.
Input: A point $f^{\ast}\in\mathcal{F}$, norm $L_{2}(\mathbb{P}_{X})$, step count $T$
1 $k\leftarrow 1$;
$\Upsilon\leftarrow[f^{\ast}\in\mathcal{F}]$ ;
  /* This array is needed solely in the proof and is not used by the estimator */
2 while $k\leq T-1$ do
       Take a $\frac{d}{2^{k}(C+1)}$ maximal packing set $M_{k}$ of the set $B\big{(}f^{\ast},\frac{d}{2^{k-1}}\big{)}\cap\mathcal{F}$ ;
        /* The packing sets should be constructed prior to seeing the data */
       $f^{\ast}\leftarrow\mathop{\mathrm{argmin}}_{f\in M_{k}}\sum_{i=1}^{n}(Y_{i}-f(X_{i}))^{2}$ ;
        /* Break ties by picking the point with the least lexicographic ordering. */
3       $\Upsilon$.append$(f^{\ast})$;
4       $k\leftarrow k+1$;
5
return $f^{\ast}$
Algorithm 1 Upper Bound Algorithm for Nonparametric Regression
To prove this algorithm has the desired properties, we first relate the $L_{2}(\mathbb{P}_{X})$ norm to the $L_{2}(\mathbb{P}_{n})$ norm. In particular, Lemma 2.5 shows that if $f,g\in\mathcal{F}$ are far apart but $f$ is close to $\bar{f}$ in $L_{2}(\mathbb{P}_{X})$-distances, then the same is true for $f,g,\bar{f}$ with high probability in the $L_{2}(\mathbb{P}_{n})$-distance. The proof, given in the appendix, relies on Bernstein’s concentration inequality. This result also applies when the noise is sub-Gaussian.
Lemma 2.5.
Suppose $f,g\in\mathcal{F}$ satisfy $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$ for some sufficiently large $C>0$. Suppose further that $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}<\delta^{2}$. Then $n\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}^{2}\leq 2\delta^{2}$ and $n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}\geq(C^{2}-1)\delta^{2}$ both simultaneously hold with probability at least
$1-\exp\left(-\frac{3\delta^{2}}{32B_{\mathcal{F}}^{2}}\right)-\exp\left(-\frac{3\delta^{2}}{8B_{\mathcal{F}}^{2}\cdot(3C^{2}+1)}\right).$
{toappendix}
@@ 4       $k\leftarrow k+1$;
We recall Bernstein’s inequality (bennetbernstein1962, see pages 33-34). Namely, given i.i.d. random variables $Z_{1},\dots,Z_{n}$ with $\mathbb{E}[Z_{1}]=0$, $|Z_{1}|\leq K$ almost surely, and any $t>0$, we have
$\mathbb{P}\left(\sum_{i}Z_{i}>t\right)$ $\leq\exp\left(-\frac{t^{2}/2}{n\mathbb{E}[Z_{1}^{2}]+Kt/3}\right).$
If we replace the $Z_{i}$ with $-Z_{i}$, a similar one-sided tail bound is derived, with $t>0$ still:
$\mathbb{P}\left(\sum_{i}Z_{i}<-t\right)$ $\leq\exp\left(-\frac{t^{2}/2}{n\mathbb{E}[Z_{1}^{2}]+Kt/3}\right).$
We set $Z_{i}=|f(X_{i})-g(X_{i})|^{2}-\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$ for each $i\in[n]$, which are mean 0 and i.i.d. random variables with $|Z_{1}|\leq K$ for $K=8B_{\mathcal{F}}^{2}$, so that $\sum_{i=1}^{n}Z_{i}=n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}.$ Moreover,
$\mathbb{E}[Z_{1}^{2}]$ $\leq\mathbb{E}|f(X_{i})-g(X_{i})|^{4}\leq 4B_{\mathcal{F}}^{2}\mathbb{E}|f(X_{i})-g(X_{i})|^{2}=K\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}.$
We set $t=C^{-2}\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$. The first and last inequality use $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$, and the second inequality is due to Bernstein.
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$ $=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot C^{2}\delta^{2}\right)$

### v5

@@ Appendix A Proofs for Section 2
We use [Wainwright, 2019, (15.52)], the decoupling property of Kullback-Leibler (KL) divergence $D_{\mathrm{KL}}$, and the decomposition of the joint distribution of $(X_{i},Y_{i})$. Let $\mathbb{P}_{Y_{i}|X_{i}}^{f_{j}}$ be the distribution of $(Y_{i}|X_{i},J=j)$, which is a normal distribution with mean $f_{j}(X_{i})$ and variance $\sigma^{2}$. Denote by $\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}$ the distribution of $(Y_{i},X_{i}|J=j)$ and $\mathbb{P}_{X_{i}}$ that of $X_{i}$, so that $\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}=\mathbb{P}_{X_{i}}\times\mathbb{P}_{Y_{i}|X_{i}}^{f_{j}}$. Note that the $X_{i}$ does not depend on $f_{j}$. Similarly, for any function $h\in\mathcal{F}$, if we set $Y_{i}=h(X_{i})+\xi_{i}$ instead, we denote the analogous distributions by $\mathbb{P}_{(X_{i},Y_{i})}^{h}=\mathbb{P}_{X_{i}}\times\mathbb{P}_{Y_{i}|X_{i}}^{h}$. Then we obtain
$I(\vec{X},\vec{Y};J)$ $\leq\frac{1}{m}\sum_{j}D_{\mathrm{KL}}\left(\prod_{i}\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}\parallel\prod_{i}\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)$
$=\frac{1}{m}\sum_{j}\sum_{i}D_{\mathrm{KL}}\left(\mathbb{P}_{(X_{i},Y_{i})}^{f_{j}}\parallel\mathbb{P}_{(X_{i},Y_{i})}^{h}\right)$
@@
$=\frac{1}{m}\sum_{j}\sum_{i}\mathbb{E}_{X_{i}}\left[D_{\mathrm{KL}}\left(\mathcal{N}(f_{j}(X_{i}),\sigma^{2})\parallel\mathcal{N}(h(X_{i}),\sigma^{2})\right)\right]$
@@
Pick a $f\in\mathcal{F}$ that achieves
$M:=\mathcal{M}_{\mathcal{F}}^{\operatorname{loc}}(\varepsilon,c):=\sup_{f\in\mathcal{F}}\mathcal{M}(\varepsilon/c,B(f,\varepsilon)\cap\mathcal{F})$
(otherwise, repeat the following argument by picking a limiting sequence of functions). Let $f_{1},\dots,f_{M}$ be a maximal $\varepsilon/c$-packing set for $B(f,\varepsilon)\cap\mathcal{F}$, so that $\|f_{i}-f_{j}\|_{L_{2}(\mathbb{P}_{X})}>\varepsilon/c$ for all $i\neq j$. Because $f_{1},\dots,f_{M}\in B(f,\varepsilon)\cap\mathcal{F}$, it follows for each $j\in[M]$ that $\|f_{j}-f\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon$. Letting $J$ be uniformly distributed on $[M]$, we have from Lemma 2.2 that
@@
$\inf_{\hat{f}}\sup_{\bar{f}\in\mathcal{F}}\mathbb{E}_{\vec{X},\vec{Y}}\|\hat{f}(\vec{X},\vec{Y})-\bar{f}\|_{L_{2}(\mathbb{P}_{X})}^{2}$ $\geq\frac{\varepsilon^{2}}{4c^{2}}\bigg{(}1-\frac{\frac{n\varepsilon^{2}}{2\sigma^{2}}+\log 2}{\log M}\bigg{)}.$
@@
Appendix B Proofs for Section 3
Remark B.1 (Breaking Ties in Algorithm).
@@ Appendix B Proofs for Section 3
By Theorem A.1 in Györfi [2002], the set of continuous functions is dense in $\mathcal{F}$. Hence we can construct all of our packing sets using continuous functions only. That is, if $\mathcal{F}_{c}$ denote the set of all continuous functions in $\mathcal{F}$, then we construct maximal packing sets $M_{k}$ of sets $B(\cdot,d/2^{k-1})\cap\mathcal{F}_{c}$ prior to pruning. This $M_{k}$ clearly has cardinality that is smaller than or equal to the $M_{k}$ constructed with $\mathcal{F}$ instead since $\mathcal{F}_{c}\subseteq\mathcal{F}$. In addition, since the set of all continuous functions is dense in $\mathcal{F}$, for any $g\in\mathcal{F}$ and $\varepsilon>0$, there exists an $h\in\mathcal{F}_{c}$ such that $\|g-h\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon$. Pick the point $f\in M_{k}$ which is $d/2^{k-1}$ close to $h$ and conclude that $\|f-g\|_{L_{2}(\mathbb{P}_{X})}\leq\varepsilon+d/2^{k-1}$. Let $\varepsilon$ go to zero, and we conclude that any point in $\mathcal{F}$ can be approximated by a point in $\mathcal{F}_{c}$ within distance $d/2^{k-1}$. Now suppose that the domain of the functions is $\mathbb{R}^{p}$. Take the set $\mathbb{Q}^{p}$ which is dense in $\mathbb{R}^{p}$. This is a well-ordered set since there is a bijection between $\mathbb{Q}^{p}$ and $\mathbb{N}$. Hence the functions in $M_{k}$ can be restricted to $\mathbb{Q}^{p}$ and then ordered lexicographically. There will not be any ties, since the functions are separated and continuous (hence the functions cannot be the same on $\mathbb{Q}^{p}$ otherwise they are the same on $\mathbb{R}^{p}$ which is a contradiction) so we can take the smallest function in the lexicographic ordering.
Proof of Proposition 3.4.
A uniform sup-norm bound trivially implies (1.1), so we just consider the $L$-sub-Gaussian case. Also note that by Stirling’s approximation [Robbins, 1955], we have
$p^{p}$ $\leq\frac{p!\exp(p)\exp\left(-\tfrac{1}{12p+1}\right)}{\sqrt{2\pi p}}=\frac{p!\cdot\exp\left(p-\frac{1}{12p+1}-\frac{1}{2}\log p\right)}{\sqrt{2\pi}}$
$=\frac{p!}{\sqrt{2\pi}}\cdot\bigg{[}\exp\underbrace{\left(1-\tfrac{1}{p(12p+1)}-\tfrac{1}{2p}\log p\right)}_{\leq 1}\bigg{]}^{p}\leq\tilde{C}e^{p}p!.$
Therefore
$\|f-g\|_{L_{2p}(\mathbb{P}_{X})}^{2p}$ $\leq\left(L\sqrt{2p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}\right)^{2p}$
$=(2L^{2})^{p}p^{p}\left(\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right)^{p}$
$=(2L^{2})^{p}p^{p}\left(\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\right)^{p-2}\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\cdot$
$\leq(2L^{2})^{p}d^{p-2}p^{p}\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$\leq(2L^{2})^{p}d^{p-2}\tilde{C}e^{p}p!\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$
$=\frac{\tilde{L}^{p-1}p!\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}}{2},$
where $\tilde{L}=(2\tilde{C})^{\tfrac{1}{p-1}}\cdot(2L^{2}e)^{\tfrac{p}{p-1}}d^{\tfrac{p-2}{p-1}}$. Note that $\tfrac{1}{p-1}\leq 1$, $\tfrac{p}{p-1}\leq 2$, and $\tfrac{p-2}{p-1}\leq 1$. Hence
$\tilde{L}\leq(2\tilde{C})(2L^{2}e)^{2}d=8e^{2}\tilde{C}L^{4}d.$
Thus, setting $L^{\prime}=8e^{2}\tilde{C}L^{4}d$, we have $\|f-g\|_{L_{2p}(\mathbb{P}_{X})}^{2p}\leq\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}{L^{\prime}}^{p-1}p!/2$ for some fixed constant $L^{\prime}>0$ and all $p\geq 2$. ∎
Proof of Example 1.1.
Observe that any difference $h=f-g$ of two functions $f,g$ in $\mathcal{F}$ satisfies $h(0)=0$ and is $2$-Lipschitz. We need to argue that for some $L$, for all $p\geq 2$ we have
$\mathbb{E}h^{2p}(X)\leq\mathbb{E}h^{2}(X)L^{p-1}p!/2.$ (B.1)
By Rademacher’s theorem, $h$ is differentiable almost everywhere, and since $h$ is $2$-Lipschitz we have $|h^{\prime}(x)|\leq 2$ at any point of differentiability $x$. In addition, $|h(x)|\leq 2|x|$. Hence by Stein’s Lemma applied to the function $\phi(x)=h^{2p}(x)/x$, we have
$\mathbb{E}h^{2p}(X)$ $=\mathbb{E}\left[\frac{h^{2p}(X)\cdot X}{X}\right]=2p\cdot\mathbb{E}\left[\frac{h^{2p-1}(X)h^{\prime}(X)}{X}\right]-\mathbb{E}\left[\frac{h^{2p}(X)}{X^{2}}\right]$
$\leq 8p\mathbb{E}\left[\frac{|h^{2p-1}(X)|}{2|X|}\right]\leq 8p\mathbb{E}[h^{2p-2}(X)].$
Repeating this logic recursively $p-2$ times,
$\mathbb{E}h^{2p}(X)\leq 8^{p-1}p!\mathbb{E}h^{2}(X).$
Thus (B.1) holds with $L=2^{(3p-2)/(p-1)}\leq 2^{4}=16$ since $\frac{3p-2}{p-1}\leq 4$ for any $p\geq 2$. Finally we argue that the class $\mathcal{F}$ has a bounded $L_{2}(\mathbb{P}_{X})$ diameter. This is so since $\mathbb{E}h^{2}(X)\leq 4\mathbb{E}X^{2}=4.$
However, this function class is clearly not bounded in sup-norm, and we can also show it is not $L$-sub-Gaussian. To see the latter, we follow the logic of the counter-examples in Han and Wellner [2019], specifically Proposition 3 where the authors show a small-ball condition defined in their Section 5.2.1 fails. [^2] The proposition and section referred to correspond to an earlier version of a preprint for this work, namely arXiv:1706.02410v1. By the Paley-Zygmund inequality, this will also imply the failure of $L$-sub-Gaussianity. Define the sequence of functions $\{f_{n}\}_{n\in\mathbb{N}}$ as follows. Let $f_{n}(x)$ be 00 everywhere except let its graph form a triangle with vertices at $(n,0)$, $(n+2,0)$, and $(n+1,1)$, or more formally,
$f_{n}(x)=\begin{cases}0&x<n\text{ or }x\geq n+2\\ x-n&x\in[n,n+1)\\ n+2-x&x\in[n+1,n+2).\end{cases}$
This clearly satisfies $f_{n}(0)=0$ and the $1$-Lipschitz requirement (hence it can be represented as a difference between a $1$-Lipschitz function and the zero function (which is also $1$-Lipschitz)). Now pick any $\eta>0$. Observe that
$\mathbb{P}(|f_{n}(X)|\geq\eta\|f_{n}\|_{L_{2}(\mathbb{P}_{X})})\leq\mathbb{P}(|f_{n}(X)|>0)\leq\mathbb{P}(X\in(n,n+2))\to 0,$
the last step using that $X\sim\mathcal{N}(0,1)$. This implies the small-ball condition fails. Hence $\mathcal{F}$ is not $L$-sub-Gaussian. ∎
@@
For $\theta>0$, we multiply the below inequality inside the probability by $\theta$ and exponentiate. Then an application of Markov’s inequality yields
$\mathbb{P}\bigg{(}\sum_{i\in[n]}(Z_{i}^{2}-\mathbb{E}Z^{2})\geq t\bigg{)}$ $\leq\frac{1}{\exp(\theta t)}\cdot\mathbb{E}\exp\bigg{(}\theta\sum_{i\in[n]}(Z_{i}^{2}-\mathbb{E}Z^{2})\bigg{)}$
$=\frac{1}{\exp(\theta t)}\cdot\prod_{i\in[n]}\mathbb{E}\exp\left(\theta(Z_{i}^{2}-\mathbb{E}Z^{2})\right).$ (B.2)
Similarly, still assuming $\theta>0$ but this time multiplying by $-\theta$ and exponentiating, we have
$\mathbb{P}\bigg{(}\sum_{i\in[n]}(Z_{i}^{2}-\mathbb{E}Z^{2}\bigg{)}\leq-t)$ $\leq\frac{1}{\exp(\theta t)}\cdot\mathbb{E}\exp\bigg{(}-\theta\sum_{i\in[n]}(Z_{i}^{2}-\mathbb{E}Z^{2})\bigg{)}$
$=\frac{1}{\exp(\theta t)}\cdot\prod_{i\in[n]}\mathbb{E}\exp\big{(}-\theta(Z_{i}^{2}-\mathbb{E}Z^{2})\big{)}.$
Now we control each $\mathbb{E}\exp(\eta\theta(Z_{i}^{2}-\mathbb{E}Z^{2}))$ from above, where $\eta\in\{\pm 1\}$. We have
$\mathbb{E}\exp(\eta\theta(Z_{i}^{2}-\mathbb{E}Z^{2}))$ $=1+\sum_{i=2}^{\infty}\frac{\theta^{i}\mathbb{E}(\eta(Z^{2}-\mathbb{E}Z^{2}))^{i}}{i!}$
$\leq 1+\sum_{i=2}^{\infty}\frac{\theta^{i}\mathbb{E}(|Z^{2}-\mathbb{E}Z^{2}|)^{i}}{i!}$
$\leq 1+\sum_{i=2}^{\infty}\frac{\theta^{i}\mathbb{E}(|Z^{2}|+|\mathbb{E}Z^{2}|)^{i}}{i!}$
$\leq 1+\sum_{i=2}^{\infty}\frac{\theta^{i}2^{i-1}(\mathbb{E}Z^{2i}+(\mathbb{E}Z^{2})^{i})}{i!}$
$\leq 1+\sum_{i=2}^{\infty}\frac{\theta^{i}2^{i}\mathbb{E}Z^{2i}}{i!},$
where we used $(\mathbb{E}Z^{2})^{i}\leq\mathbb{E}Z^{2i}$ by Jensen’s inequality, and $(a+b)^{i}\leq 2^{i-1}(a^{i}+b^{i})$ for $a,b>0$ (which follows from Jensen’s inequality or induction). Now applying our hypothesis on the $2i$-th moment and recognizing a geometric sum, we have
$\mathbb{E}\exp(\theta\eta(Z_{i}^{2}-\mathbb{E}Z^{2}))$ $\leq 1+\sum_{i=2}^{\infty}\frac{(2\theta L)^{i}\mathbb{E}Z^{2}}{2L}$
$=1+(2\theta L)^{2}\mathbb{E}Z^{2}\sum_{i=0}^{\infty}\frac{(2\theta L)^{i}}{2L}$
$=1+\frac{(2\theta)^{2}\mathbb{E}Z^{2}L}{2}\frac{1}{1-2\theta L}$
$\leq\exp\left((2\theta)^{2}\mathbb{E}Z^{2}L\cdot\frac{1}{1-2\theta L}\right),$ (B.3)
provided that $12\theta L<$ 1. The final inequality used the result $1+x\leq e^{x}$ for all $x\in\mathbb{R}$.
Next, select $\theta=\frac{t}{8n\mathbb{E}Z^{2}L+2Lt}$, so that we indeed have $12L\theta<$ 1. Thus $\tfrac{1}{1-2\theta L}=\frac{8n\mathbb{E}Z^{2}L+2Lt}{8n\mathbb{E}Z^{2}L}$. Combining (B.2) and (B.3),
$\mathbb{P}(\sum(Z_{i}^{2}-\mathbb{E}Z^{2})\geq t)$
$\leq\exp\left((2\theta)^{2}\mathbb{E}Z^{2}L\cdot\frac{1}{1-2\theta L}-\theta t\right)$
$=\exp\left(\frac{(2t)^{2}\mathbb{E}Z^{2}L}{(8n\mathbb{E}Z^{2}L+2Lt)^{2}}\frac{8n\mathbb{E}Z^{2}L+2Lt}{8n\mathbb{E}Z^{2}L}-\frac{t^{2}}{8n\mathbb{E}Z^{2}L+2Lt}\right)$
$=\exp\left(\frac{4t^{2}}{8n\cdot(8n\mathbb{E}Z^{2}L+2Lt)}-\frac{t^{2}}{8n\mathbb{E}Z^{2}L+2Lt}\right)$
$\leq\exp\left(\frac{4t^{2}}{8\cdot(8n\mathbb{E}Z^{2}L+2Lt)}-\frac{t^{2}}{8n\mathbb{E}Z^{2}L+2Lt}\right)$
$=\exp\left(-\frac{t^{2}}{16n\mathbb{E}Z^{2}L+4Lt}\right),$
which is what we aimed to show. Finally note that the same bound applies for the other tail. ∎
Proof of Lemma 3.7.
Suppose we take $Z_{i}=|f(X_{i})-g(X_{i})|$ for each $i\in[n]$, noting these are i.i.d. and $\mathbb{E}Z_{i}^{2p}\leq\mathbb{E}Z_{i}^{2}L^{p-1}p!/2$ by (1.1). Note that $\sum_{i=1}^{n}Z_{i}^{2}=\sum_{i=1}^{n}|f(X_{i})-g(X_{i})|^{2}=n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}$ and $\mathbb{E}[Z_{1}^{2}]=\|f-g\|_{L_{2}(\mathbb{P}_{n})}$. Then by Lemma 3.6
$\mathbb{P}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq t\right)$
$\vee\mathbb{P}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}-n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\leq-t\right)$
$\leq\exp\left(-\frac{t^{2}}{16nL\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}+4Lt}\right).$ (B.4)
An identical bound holds for $\|f-\bar{f}\|_{L_{2}(\mathbb{P}_{n})}$.
Set $t=C^{-2}\cdot n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}$. Then we derive the following, where the first and last inequality use $n\|f-g\|_{L_{2}(\mathbb{P}_{X})}^{2}\geq C^{2}\delta^{2}$, and the second inequality uses (B.4).
$\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(C^{2}-1)\delta^{2}\right)$
$=\mathbb{P}_{\bar{f}}\left(n\|f-g\|_{L_{2}(\mathbb{P}_{n})}^{2}<(1-C^{-2})\cdot C^{2}\delta^{2}\right)$